OpenStudy (anonymous):

Add and answer in simplest fractional form. -4/3 + 7/8 =

OpenStudy (anonymous):

-11/24

OpenStudy (lalaly):

\[\frac{-32}{24} + \frac{21}{24} = \frac{-11}{24}\]

OpenStudy (anonymous):

yes, -11/24:)